OpenStudy (anonymous):

Find the inverse S(h)=1000+10h

OpenStudy (saifoo.khan):

(h-1000) / 10 = h

OpenStudy (anonymous):

y = 1000 + 10h y - 1000 = 10h h = (y - 1000) / 10 g(y) = (y = 1000) / 10 where g(y) is the inverse